OpenStudy (anonymous):

Perform the operation (x-y/2)-(x+y/3)

OpenStudy (anonymous):

(x-y/2)-(x+y/3) = (x-3y/6)-(x+2y/6) = x-3y/6-x-2y/6 = y/6

OpenStudy (anonymous):

\[\frac{x-y}{2}-\frac{x+y}{3}=\frac{3(x-y)-2(x+y)}{2*3}\]

OpenStudy (anonymous):

\[\frac{(3x-3y)-(2x+2y)}{6}=\frac{3x-3y-2x-2y}{6}\]

OpenStudy (anonymous):

\[\frac{x-5y}{6}\]

OpenStudy (anonymous):

The question is ambiguous as both answers are correct depending on the question